OpenStudy (jordan2231):

When Clark was 7 years old he had 22 rare coins. Each year Clark added 15 coins to his collection. Assuming Clark does not sell or lose any coins, how old will Clark be when he has 577 coins? A) 30 B) 31 C) 37 D) 44

OpenStudy (jordan2231):

@Elson

OpenStudy (elson):

Take 577-22=555.555/15=37.37+7=44 44 yrs old or D